A flagpole that is 35 feet high casts a 50 ft shadow. At the time of day, how long would the shadow be of a girl who is 5’9?
Thanks!

Found 3 solutions by josgarithmetic, Boreal, Alan3354:
Answer by josgarithmetic(39617)
Answer by Boreal(15235)

5'9=69 inches
35/50=69/x
35x=3450
x=98.57 inches or 8 feet 1.57 inches.

Answer by Alan3354(69443)

A flagpole that is 35 feet high casts a 50 ft shadow. At the time of day, how long would the shadow be of a girl who is 5’9?
--------------
5'9" = 5.75 feet
============
35/50 = 5.75/shadow